OpenStudy (torii897):

please help!! Calculate the area of the regular pentagon below: A regular pentagon with side length of 11.8 inches and dotted line from center to middle of side of 8.1 inches.

OpenStudy (mrnood):

it is simply 5 triangle base = 11.8, height = 8.1 area of triangle - 0.5 base * height

OpenStudy (torii897):

and I multiply that

OpenStudy (jack1):

\(\Large A = \frac 14 \times n \times a^2 \times cot \frac {pi}{n} \) A | area n | number of sides a | side length

OpenStudy (jack1):

pentagon has 5 sides... so n=5 and side length is 11.8... so a = 11.8